OpenStudy (anonymous):

MEDAL ALERT! Choose the equation below that represents the line passing through the point (1, -4) with a slope of one half. y - 4 = one half(x - 1) y + 4 = 2(x - 1) y + 4 = one half(x - 1) y - 4 = 2(x + 1)

OpenStudy (anonymous):

use this formula : y - y1 = m(x - x1) (1,-4) x1 = 1 and y1 = -4 slope(m) = 1/2 now just sub your info into the formula y - (-4) = 1/2(x - 1) y + 4 = 1/2(x - 1) --> your answer any questions ?

OpenStudy (anonymous):

So, what is the base formula I'd need to remember for other similar questions?

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (gorv):

y-y1=m(x-x1) m= slope and (x1,y1) is the point through which line is passing

OpenStudy (anonymous):

Thanks a bunch!